OpenStudy (anonymous):

Need help with polynomials... (a + b)(a2 - ab + b2) a^3+ab2^+ab2-b3 (a - b)(a2 + ab + b2) a3+ab2 I dont think im doing any of this right all the (a's,b's) really throw me off

OpenStudy (johnweldon1993):

(a + b)(a^2 - ab + b^2) Alright what you do...is take the first term in the first set of parenthesis * 'a' in this case* and multiply it by every term in the second set of parenthesis....then do the same for the second term in the first set ...'b' so we have a * a^2 = a^3 a * -ab = -a^2b a * b^2 = ab^2 then b * a^2 = a^2b b * -ab = -ab^2 b * b^2 = b^3 So altogether you have a^3 - a^2b + ab^2 + a^2b - ab^2 + b^3 ^ ^ ^ ^ cancel a^3 + b^3 is your final answer for that first one...

OpenStudy (anonymous):

can u help me

OpenStudy (anonymous):

with what

OpenStudy (anonymous):

my problem

OpenStudy (anonymous):

what problem

OpenStudy (anonymous):

openstudy.com/study#/updates/52781fa3e4b0ebf8902964ff